OpenStudy (anonymous):

11v-44 over6 divided by v-4 over 33v

OpenStudy (anonymous):

\[\frac{11v-44}{6} \div\frac{v-4}{33v}\] \[=11\frac{v-4}{6} * \frac{33v}{v-4}\] \[=\frac{11*11*v}{2}=\frac{121v}{2}\]